OpenStudy (anonymous):

Solve for x. 4x – (6 – x) = 12 + 4x A. –18 B. 2 C. –2 D. 18

OpenStudy (anonymous):

What is your guess?

OpenStudy (anonymous):

D 18

OpenStudy (anonymous):

4x – (6 – x) = 12 + 4x -4x -4x -6+x = 12 +6 +6 x= 18

OpenStudy (anonymous):

Dragonquill, explain please. Coc